# Evidence for large baryonic feedback at low and intermediate redshifts from kinematic Sunyaev-Zel’dovich observations with ACT and DESI photometric galaxies

### v1

@@ Acknowledgements.
Appendix B Effect of removing corrections and comparison with Data Release 9
@@
To corroborate the claim that the additional cleaning applied in this work has minimal effect on the profile shapes, in Fig. 8, we present a comparison between the fiducial (cleaned) analysis and the case where neither of these corrections is applied. Visually, the curves bear a very strong resemblance with each other, all measurements being within 1 of each other. Most prominently, we see that the third bin behaves more physically in the corrected case (since we adopt a CAP filter, we expect the measured profiles to increase until all gas is encompassed within some aperture radius), while the signal in the first bin gets a slight boost when no corrections are applied. This is reassuring to see, as it suggests that our cleaning procedure works as expected (i.e., it does not significantly bias our outputs and reduces the anomalous behavior exhibited by the third bin).

### v2

@@ Appendix F Comparison with CMASS measurement
Appendix G Effect of removing corrections and comparison with Data Release 9
@@
To corroborate the claim that the additional cleaning applied in this work has minimal effect on the profile shapes, in Fig. 9, we present a comparison between the fiducial (cleaned) analysis and the case where neither of these corrections is applied. Visually, the curves bear a very strong resemblance with each other, all measurements being within 1 of each other. Most prominently, we see that the third bin behaves more physically in the corrected case (since we adopt a CAP filter, we expect the measured profiles to increase until all gas is encompassed within some aperture radius777In particular, in the unphysical case of an isolated halo, the profile would flatten out once all the gas is encompassed. However, due to the presence of other correlated structure nearby (i.e., two-halo term), we expect the profiles to keep increasing mildly with radius. In reality, on large scales, the noise is highly correlated and dominated by the large variance of the primary CMB.), while the signal in the first bin gets a slight boost when no corrections are applied. This is reassuring to see, as it suggests that our cleaning procedure works as expected (i.e., it does not significantly bias our outputs and reduces the anomalous behavior exhibited by the third bin).